OpenStudy (anonymous):

Find the angle between u=-4i+3j and w=-5i-7j

OpenStudy (anonymous):

can you find the angle for the individual vector????

OpenStudy (anonymous):

think about dot product

OpenStudy (anonymous):

add like times -4j+3j is u=-j and w=-12j so it has to be between -1j and -12j

OpenStudy (anonymous):

steps: first find the angle of U then find the angle of W subtract the smaller angle from the larger, the result is the angle between these vectors.

terenzreignz (terenzreignz):

Since dot product is also given as... \[\Large \vec u \cdot \vec w = |\vec u||\vec w|\cos(\theta)\] Then it follows that \[\Large \cos(\theta) = \frac{\vec u \cdot \vec w}{|\vec u||\vec w|}\]